OpenStudy (epikrebel):

Solve for x: x^2+1=0. please explain.

OpenStudy (anonymous):

\[x^2=-1\] \[x=\pm\sqrt{-1}\]

OpenStudy (anonymous):

or if you prefer \[x=\pm i\]

OpenStudy (campbell_st):

this will have complex number answer subtract 1 from both sides \[x^2 = -1\] take the square root of both sides let \[-1 = i^2\] so you have \[x = \pm \sqrt{i^2}\] I'll leave you to finish the problem.

OpenStudy (epikrebel):

Okay I get it now. Thanks a lot.